OpenStudy (anonymous):

2x+2=x+1 solve for x

OpenStudy (anonymous):

i tried but i failed it is very hard

OpenStudy (anonymous):

foil

OpenStudy (angela210793):

i think u have to divide 2 both sides so u can find y

OpenStudy (anonymous):

yeah

OpenStudy (akshay_budhkar):

you can integrate to get your answer as -1

OpenStudy (angela210793):

Ishaan agreed with me.....---> I'm smart ^_^

OpenStudy (akshay_budhkar):

or differentiate

OpenStudy (anonymous):

hahahahaha, whaat? XD

OpenStudy (anonymous):

2x-x = 1-2 x=-1

OpenStudy (angela210793):

no i think sqr both sides would b better

OpenStudy (anonymous):

:X

OpenStudy (anonymous):

oh that is wrong musifus

OpenStudy (angela210793):

soooooooooooooo wrong....:O

OpenStudy (anonymous):

._. why?

OpenStudy (anonymous):

you can't do that because you are lifesaver you have to obey seniors than you

OpenStudy (anonymous):

hahahahahahahahahahahha

OpenStudy (anonymous):

what they are telling is correct

OpenStudy (angela210793):

hahahhahhahahhaah :D yea

OpenStudy (anonymous):

you are doing it very wrong

OpenStudy (anonymous):

but this guy needs help XD, dont confuse him more than he already is

OpenStudy (anonymous):

lol what makes you thing he needs help

OpenStudy (anonymous):

he made a question

OpenStudy (angela210793):

naive :P

OpenStudy (anonymous):

:o?

OpenStudy (anonymous):

okay then go to my profile I have some confusions there will you do that please my profile and then questions asked by me

OpenStudy (angela210793):

jkjkjk :D :D